# Interpretability of linear regression models of glassy dynamics

### v1

@@ Appendix B Ordinary least squares regression and Ridge regression
Appendix C Condition number of a matrix
We briefly review the condition number of a matrix, without going into the details of a mathematically rigorous treatment. It is defined using the norm of a matrix, and we also review its meaning as a measure of a matrix’s instability. One can see detailed discussions in, e.g., Refs. 79, 80.
C.1 Matrix norm
@@
C.2 Definition of condition number
@@ C.3 Instability of matrix
Appendix D Ridge regression in the PCA basis
D.1 Expression of the Pearson coefficient
@@
D.2 Role of the regularization parameter
@@
It is then easy to find the Pearson coefficient, see the Appendix D,
@@
D.3 Suppression of the oscillatory behavior of the weights

### v2

@@ Appendix C Ordinary least squares regression and Ridge regression
Appendix D Condition number of a matrix
We briefly review the condition number of a matrix, without going into the details of a mathematically rigorous treatment. It is defined using the norm of a matrix, and we also review its meaning as a measure of a matrix’s instability. One can see detailed discussions in, e.g., Refs. 81, 82.
D.1 Matrix norm
@@
D.2 Definition of condition number
@@ D.3 Instability of matrix
Appendix E Ridge regression in the PCA basis
E.1 Expression of the Pearson coefficient
@@
E.2 Role of the regularization parameter
@@
It is then easy to find the Pearson coefficient
@@
E.3 Suppression of the oscillatory behavior of the weights